OpenStudy (anonymous):

While preparing to visit the United States, Greg exchanged 200 Canadian dollars for 1/5 fewer U.S dollars. After spending ¾ of the U.S dollars, how many did he have left?

OpenStudy (anonymous):

(1/5) 200 = 40 200 - 40 = 160 US$ (1/4) 160 = 40 US$